OpenStudy (anonymous):

If f(x)=5x-3 then what is f^1(x)?

Parth (parthkohli):

Do you mean \(f^{-1}(x)\)?

Parth (parthkohli):

You first multiply 5 to a number, then you subtract 3. You'd do the exactly opposite thing. You first add 3, then you divide 3 from the whole.

Parth (parthkohli):

There's a technical way to do that too.

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(x) = 5x - 3 \Longrightarrow y = 5x - 3}\) \( \color{Black}{\Rightarrow y + 3 = 5x}\) \( \color{Black}{\Rightarrow \Large {y + 3 \over 5} \normalsize = x}\)

Parth (parthkohli):

As you have solved for \(x \) in terms of \(y\), you may now replace \(x\) with \(f^{-1}(x)\) and \(y \) with \(x\).

Parth (parthkohli):

That gets you \[ f^{-1}(x) = {x + 3 \over 5}\]